# Many (most?) column subset selection criteria

### v1

@@ Abstract
We consider a variety of criteria for selecting $k$ representative columns from a real matrix $\boldsymbol{A}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{A})\geq k$. The criteria include the following optimization problems: absolute volume and S-optimality maximization; norm and condition minimization in the two-norm , Frobenius norm and Schatten $p$-norms for $p>2$; stable rank maximization; and the new criterion of relative volume maximization. We show that these criteria are NP hard and do not admit polynomial time approximation schemes (PTAS). To formulate the optimization problems as decision problems, we derive optimal values for the subset selection criteria, as well as expressions for partitioned pseudo-inverses.
@@ 1 Introduction
Given a general [^2] We make no assumptions on the dimensions $m$ and $n$ of $\boldsymbol{A}$. matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{A})\geq k$, what does it mean to find $k$ representative columns $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ of $\boldsymbol{A}$? This problem is known as ‘column subset selection.’ Our goal is to establish the complexity classes for the popular subset selection problems in Table 1.
@@
Having established that a criterion is NP hard, it is reasonable to ask if it can be approximated in polynomial time. This leads to the concept of a Polynomial Time Approximation Scheme (PTAS). A PTAS is a polynomial time algorithm such that for each $\epsilon>0$ there exists a $(1+\epsilon)$-approximation algorithm for minimization problems, and a $(1-\epsilon)$-approximation algorithm for maximization problems [31, Section 1], [34, Definition 1.2]. We show that most criteria in Table 1 do not admit a PTAS either.
@@ 1.1 Complexity Theory
We focus on decision problems with a binary output ‘yes’ or ‘no’, and adopt the Turing model of computation, where inputs are rational numbers and the input size is measured in the number of bits.
@@
The complexity class $P$ consists of decision problems that a deterministic Turing machine can solve in polynomial time, while the complexity class $NP$ consists of decision problems for which a deterministic Turing machine can verify an output of ‘yes’ in polynomial time. A problem is NP hard if there is polynomial time reduction to this problem from any problem in NP. A decision problem is NP complete if it is in NP and it is also NP hard. Thus, an NP hard problem is at least as hard as an NP complete problem [17, Section 5.1].
@@
A standard strategy for establishing NP hardness performs a polynomial time reduction from an NP complete problem to the problem at hand. This reduction makes it possible to convert any algorithm for solving the problem at hand into a corresponding algorithm for an NP complete problem. In particular, a reduction that operates in polynomial time makes it possible to convert any polynomial time for the problem at hand to a corresponding polynomial time algorithm for the NP complete problem [17, Section 1.5]. Following [9], we choose the NP complete problem ‘Exact cover by 3 sets (X3C)’ [17, Section A.3.1], which appears in Karp’s 1972 list of NP complete problems [24, Section 4].
@@
To show that a subset selection criterion does not admit a PTAS, we create a “gap” between the optimal value and a particular suboptimal value, the so-called threshold, and show that it is NP hard to approximate the criterion beyond the threshold, that is, for any value inside the gap.
@@
Unfortunately, creating such a gap is not always feasible [31, Section 1], and we were not able to do so for a few criteria, including Schatten $p$-norm minimization for $p>2$ (Section 4) and condition number minimization in Schatten $p$-norms for $p>2$ (Section 6).
@@
Quantity Criterion NP hard No PTAS Definition Volume $\max_{\boldsymbol{C}}{\operatorname{\mathrm{vol}}(\boldsymbol{C})}$ [9, Theorem 4] [9, Theorem 6](1) $\operatorname{\mathrm{vol}}(\boldsymbol{C})\equiv\prod_{j=1}^{k}{\sigma_{j}(\boldsymbol{C})}$ Theorem 3 Theorem 4 $\max_{\boldsymbol{C}}{\operatorname{\mathrm{rvol}}(\boldsymbol{C})}$ Theorem 1 Theorem 2 $\operatorname{\mathrm{rvol}}(\boldsymbol{C})\equiv\operatorname{\mathrm{vol}}(\boldsymbol{C})/\|\boldsymbol{C}\|_{2}^{k}$ $\max_{\boldsymbol{C}}{\operatorname{\mathrm{s_{opt}}}(\boldsymbol{C})}$ Theorem 3 Theorem 5 $\operatorname{\mathrm{s_{opt}}}(\boldsymbol{C})\equiv(\operatorname{\mathrm{vol}}(\boldsymbol{C})/\prod_{i=1}^{k}{\|\boldsymbol{C}\boldsymbol{e}_{i}\|_{2})^{1/k}}$ Norm $\min_{\boldsymbol{C}}{\|\boldsymbol{C}\|_{2}}$ [9, Theorem 4] [9, Theorem 5] Theorem 6 Theorem 7 $\min_{\boldsymbol{C}}{\|\boldsymbol{C}\|_{F}}$ No N/A $\min_{\boldsymbol{C}}{\|\boldsymbol{C}\|_{(p)}}$ Theorem 6 Pseudo-inverse $\min_{\boldsymbol{C}}{\|\boldsymbol{C}^{\dagger}\|_{2}}$ [9, Theorem 4] [9, Theorem 7] Theorem 8 Theorem 9 $\min_{\boldsymbol{C}}{\|\boldsymbol{C}^{\dagger}\|_{F}}$ Theorem 8 Theorem 10 $\min_{\boldsymbol{C}}{\|\boldsymbol{C}^{\dagger}\|_{(p)}}$ Theorem 8 Theorem 11 Condition no $\min_{\boldsymbol{C}}{\kappa_{2}(\boldsymbol{C})}$ [9, Theorem 4] [9, Theorem 8] $\kappa_{2}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{2}\|\boldsymbol{C}^{\dagger}\|_{2}$ Theorem 12 Theorem 13 $\min_{\boldsymbol{C}}{\kappa_{F}(\boldsymbol{C})}$ Theorem 12 Theorem 14 $\kappa_{F}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{F}\|\boldsymbol{C}^{\dagger}\|_{F}$ $\min_{\boldsymbol{C}}{\kappa_{D}(\boldsymbol{C})}$ Theorem 12 Theorem 15 $\kappa_{D}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{F}\|\boldsymbol{C}^{\dagger}\|_{2}$ $\min_{\boldsymbol{C}}{\kappa_{D,p}(\boldsymbol{C})}$ Theorem 12 $\kappa_{D,p}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{(p)}\|\boldsymbol{C}^{\dagger}\|_{2}$ Stable rank $\max_{\boldsymbol{C}}{\operatorname{\mathrm{sr}}(\boldsymbol{C})}$ Theorem 16 Theorem 17 $\operatorname{\mathrm{sr}}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{F}^{2}/\|\boldsymbol{C}\|_{2}^{2}$ $\max_{\boldsymbol{C}}{\operatorname{\mathrm{sr}}_{(p)}(\boldsymbol{C})}$ Theorem 16 $\operatorname{\mathrm{sr}}_{(p)}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{(p)}^{p}/\|\boldsymbol{C}\|_{2}^{p}$ Residual $\min_{\boldsymbol{C}}{\rho_{2}(\boldsymbol{C})}$ $\rho_{2}(\boldsymbol{C})\equiv\|(\boldsymbol{I}-\boldsymbol{C}\boldsymbol{C}^{\dagger})\boldsymbol{A}\|_{2}$ $\min_{\boldsymbol{C}}{\rho_{F}(\boldsymbol{C})}$ [33, Theorem 2.2] [6, Theorem 1.1](2) $\rho_{F}(\boldsymbol{C})\equiv\|(\boldsymbol{I}-\boldsymbol{C}\boldsymbol{C}^{\dagger})\boldsymbol{A}\|_{F}$
@@
Table 1: Complexity of criteria for selecting $k$ representative columns $\boldsymbol{C}$ from a matrix $\boldsymbol{A}$: (1) Exponential inapproximability of the volume is shown in [10, Theorem 5]. (2) Minimizing the Frobenius norm residual is NP complete [33, Theorem 2.2], and is PTAS inapproximable conditional on the Unique Games Conjecture being true [6, Theorem 1.1].
@@ 1.3 Contributions
    We introduce the criterion of ‘relative volume maximization’ and show that it is NP-hard and does not admit a PTAS (Section 2).
@@
    We derive optimal values the subset selection criteria (Section 8), so that their optimization versions can be formulated as decision problems.
@@ 1.4 Literature
We review existing work on column subset selection, focussed on computational complexity. Early surveys of column subset selection algorithms include [8, 21].
@@
NP hardness and PTAS inapproximability are shown in [9, Theorems 4 and 7] for the two-norm. Algorithms for wide matrices $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with $m<n$ are given in [1, 2, 25, 30].
@@
NP hardness and PTAS inapproximability are shown in [9, Theorems 4 and 6], while the exponential inapproximability of the volume is presented in [10, Theorem 5].
Greedy algorithms are presented in [9, 14], and local deterministic and randomized algorithms in [20, 19, 29]. This is extended to Hilbert spaces in [26]. The importance of local volume maximization for revealing the rank is demonstrated in [13].
The S-optimality criterion bears a resemblance to the relative volume and was introduced in [32, (3.5)] and [27, (3.10)].
@@
The NP completeness in the Frobenius norm is established in [33, Theorem 2.2], and PTAS inapproximability conditional on the Unique Games Conjecture being true [6, Theorem 1.1]. Algorithms for the two- and Frobenius norms are given in [5].
@@ 1.5 Assumptions
Singular value interlacing [22, 7.3.P44] implies
@@
We denote by $\boldsymbol{I}_{n}\in{\mathbb{R}}^{n\times n}=\begin{bmatrix}\boldsymbol{e}_{1}&\cdots&\boldsymbol{e}_{n}\end{bmatrix}$ the identity matrix, and by $\boldsymbol{e}_{j}\in\mathbb{R}^{n}$ its columns.

### v2

@@ Abstract
We consider a variety of criteria for selecting $k$ representative columns from a real $m\times n$ matrix $\boldsymbol{A}$, when sufficiently few columns are required, i.e., $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$. The criteria include the following optimization problems: absolute volume and S-optimality maximization; norm, pseudo-inverse norm, and condition minimization number in the two-norm, Frobenius norm and Schatten $p$-norms for $p>2$; stable rank maximization; and the new criterion of relative volume maximization, which is inversely proportional to a power of the condition number. We show that these criteria are NP hard and many do not admit polynomial time approximation schemes (PTAS). To formulate the optimization problems as decision problems, we derive optimal values for the subset selection criteria, as well as expressions for partitioned pseudo-inverses. The results for minimization of the pseudo-inverse in the Frobenius norm are applicable to trace optimization in A-optimal design.
@@ 1 Introduction
Given a general [^2] We make no assumptions on the dimensions $m$ and $n$ of $\boldsymbol{A}$. matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$, what does it mean to find $k$ representative columns $\boldsymbol{C}\in\mathbb{R}^{m\times k}$? This problem is known as ‘column subset selection’. Our goal is to establish the complexity classes for the popular subset selection problems in Table 1, in the case when sufficiently few columns are required, i.e., $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$.
@@
Having established that a criterion is NP hard, it is reasonable to ask if it can be approximated in polynomial time. This leads to the concept of a Polynomial Time Approximation Scheme (PTAS). A PTAS is a polynomial time algorithm such that for each $\epsilon>0$ there exists a $(1+\epsilon)$-approximation algorithm for minimization problems, and a $(1-\epsilon)$-approximation algorithm for maximization problems [38, Section 1], [43, Definition 1.2]. We show that most criteria in Table 1 do not admit a PTAS either.
@@ 1.1 Complexity Theory
We focus on decision problems with a binary output ‘yes’ or ‘no’, and adopt the Turing model of computation, where inputs are rational numbers, and the input size is measured in terms of bits. The only irrational number used in our reduction is $1/\sqrt{3}$, which can also be expressed in terms of a finite number of bit [24, Section 9].
@@
The complexity class $P$ consists of decision problems that can be solved by a deterministic Turing machine in polynomial time, while the complexity class $NP$ consists of decision problems for which a deterministic Turing machine can verify an output of ‘yes’ or ‘no’ in polynomial time. A problem is NP hard if there is polynomial time reduction to this problem from any problem in NP. A decision problem is NP complete if it is in NP and it is also NP hard. Thus, an NP hard problem is at least as hard as an NP complete problem [20, Section 5.1].
@@
A standard strategy for establishing NP hardness performs a polynomial time reduction from an NP complete problem to the problem at hand. This reduction makes it possible to convert any algorithm for solving the problem at hand into a corresponding algorithm for an NP complete problem. In particular, a reduction that operates in polynomial time makes it possible to convert any polynomial time for the problem at hand to a corresponding polynomial time algorithm for the NP complete problem [20, Section 1.5]. Following [9], we choose the NP complete problem ‘Exact cover by 3 sets (X3C)’ [20, Section A.3.1], which appears in Karp’s 1972 list of NP complete problems [28, Section 4].
@@
To show that a subset selection criterion does not admit a PTAS, we create a “gap” between the optimal value and a particular suboptimal value, the so-called inapproximability threshold, and show that it is NP hard to approximate the criterion beyond the threshold, that is, for any value inside the gap.
@@
Unfortunately, creating such a gap is not always feasible [38, Section 1], and we were not able to do so for a few criteria, including Schatten $p$-norm minimization for $p>2$ (Section 4) and condition number minimization in Schatten $p$-norms for $p>2$ (Section 6).
@@
Quantity Criterion NP hard No PTAS Definition Volume $\max_{\boldsymbol{C}}{\operatorname{\mathrm{vol}}(\boldsymbol{C})}$ [9, Theorem 4] [9, Theorem 6](1) $\operatorname{\mathrm{vol}}(\boldsymbol{C})\equiv\prod_{j=1}^{k}{\sigma_{j}(\boldsymbol{C})}$ [18, Theorem 8] Theorem 3 Theorem 4 $\max_{\boldsymbol{C}}{\operatorname{\mathrm{rvol}}(\boldsymbol{C})}$ Theorem 1 Theorem 2 $\operatorname{\mathrm{rvol}}(\boldsymbol{C})\equiv\operatorname{\mathrm{vol}}(\boldsymbol{C})/\|\boldsymbol{C}\|_{2}^{k}$ $\max_{\boldsymbol{C}}{\operatorname{\mathrm{s_{opt}}}(\boldsymbol{C})}$ Theorem 3 Theorem 5 $\operatorname{\mathrm{s_{opt}}}(\boldsymbol{C})\equiv(\operatorname{\mathrm{vol}}(\boldsymbol{C})/\prod_{i=1}^{k}{\|\boldsymbol{C}\boldsymbol{e}_{i}\|_{2})^{1/k}}$ Norm $\min_{\boldsymbol{C}}{\|\boldsymbol{C}\|_{2}}$ [9, Theorem 4] [9, Theorem 5] Theorem 6 Theorem 7 $\min_{\boldsymbol{C}}{\|\boldsymbol{C}\|_{F}}$ No N/A $\min_{\boldsymbol{C}}{\|\boldsymbol{C}\|_{(p)}}$ Theorem 6 Pseudo-inverse $\min_{\boldsymbol{C}}{\|\boldsymbol{C}^{\dagger}\|_{2}}$ [9, Theorem 4] [9, Theorem 7] Theorem 8 Theorem 9 $\min_{\boldsymbol{C}}{\|\boldsymbol{C}^{\dagger}\|_{F}}$ Theorem 8 Theorem 10 $\min_{\boldsymbol{C}}{\|\boldsymbol{C}^{\dagger}\|_{(p)}}$ Theorem 8 Theorem 11 Condition no $\min_{\boldsymbol{C}}{\kappa_{2}(\boldsymbol{C})}$ [9, Theorem 4] [9, Theorem 8] $\kappa_{2}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{2}\|\boldsymbol{C}^{\dagger}\|_{2}$ Theorem 12 Theorem 13 $\min_{\boldsymbol{C}}{\kappa_{F}(\boldsymbol{C})}$ Theorem 12 Theorem 14 $\kappa_{F}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{F}\|\boldsymbol{C}^{\dagger}\|_{F}$ $\min_{\boldsymbol{C}}{\kappa_{D}(\boldsymbol{C})}$ Theorem 12 Theorem 15 $\kappa_{D}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{F}\|\boldsymbol{C}^{\dagger}\|_{2}$ $\min_{\boldsymbol{C}}{\kappa_{D,p}(\boldsymbol{C})}$ Theorem 12 $\kappa_{D,p}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{(p)}\|\boldsymbol{C}^{\dagger}\|_{2}$ Stable rank $\max_{\boldsymbol{C}}{\operatorname{\mathrm{sr}}(\boldsymbol{C})}$ Theorem 16 Theorem 17 $\operatorname{\mathrm{sr}}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{F}^{2}/\|\boldsymbol{C}\|_{2}^{2}$ $\max_{\boldsymbol{C}}{\operatorname{\mathrm{sr}}_{(p)}(\boldsymbol{C})}$ Theorem 16 $\operatorname{\mathrm{sr}}_{(p)}(\boldsymbol{C})\equiv\|\boldsymbol{C}\|_{(p)}^{p}/\|\boldsymbol{C}\|_{2}^{p}$ Residual $\min_{\boldsymbol{C}}{\rho_{2}(\boldsymbol{C})}$ $\rho_{2}(\boldsymbol{C})\equiv\|(\boldsymbol{I}-\boldsymbol{C}\boldsymbol{C}^{\dagger})\boldsymbol{A}\|_{2}$ $\min_{\boldsymbol{C}}{\rho_{F}(\boldsymbol{C})}$ [40, Theorem 2.2] [6, Theorem 1.1](2) $\rho_{F}(\boldsymbol{C})\equiv\|(\boldsymbol{I}-\boldsymbol{C}\boldsymbol{C}^{\dagger})\boldsymbol{A}\|_{F}$
@@
Table 1: Complexity of criteria for selecting $k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$ representative columns $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ from $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$: (1) Exponential inapproximability of the volume is shown in [10, Theorem 5]. (2) Minimizing the Frobenius norm residual is NP complete [40, Theorem 2.2], and is PTAS inapproximable conditional on the Unique Games Conjecture being true [6, Theorem 1.1].
@@ 1.3 Contributions
    We introduce the criterion of ‘relative volume maximization’ and show that it is NP-hard and does not admit a PTAS. Unlike the volume, the relative volume is able to detect whether a matrix is ill-conditioned, because it is inversely proportional to a power of the condition number (Section 2).
@@
    We derive optimal values for the subset selection criteria (Section 8), so that their optimization versions can be formulated as decision problems.
@@
1.4 Roadmap
We give a brief sketch of our approaches for deriving NP hardness results (Section 1.4.1) and PTAS results (Section 1.4.2) for the above subset selection criteria.
1.4.1 Extension of the NP-hardness results from [9]
We perform the following steps.
  1. 1.
    Polynomial time reduction of the subset selection criteria to X3C by representing the collection of subsets as a matrix $\boldsymbol{A}$ with unit norm columns.
  2. 2.
    For such matrices with unit-norm columns, we establish optimal values for the subset selection criteria (Section 8), and show that the optimal values are achieved by matrices with orthonormal columns.
    Optimal values are derived for the criteria: maximal volume and S-optimality (Lemma 1), maximal relative volume (Lemma 2), minimal norms (Lemma 4), minimal pseudo-inverse norms (Lemma 5), minimal condition numbers (Lemma 6), and maximal stable ranks (Lemma 8).
  3. 3.
    According to [9], X3C is true is and only if $\boldsymbol{A}$ has a submatrix with orthonormal columns.
  4. 4.
    Combining the two previous items implies that the criterion is optimal if and only if X3C is true.
    This shows NP hardness of the criteria: relative volume maximization (Theorem 1), volume and S-optimality maximization (Theorem 3), norm minimization (Theorem 6), pseudo-inverse norm minimization (Theorem 8), condition number minimization (Theorem 12), and stable rank maximization (Theorem 16).
1.4.2 Systematic approach for extending PTAS results from [9]
We perform the following steps.
  1. 1.
    Monotonicity: Singular value interlacing implies that deleting columns from a matrix can only increase its relative volume (Lemma 3) and decrease the condition number (Lemma 7). In other words, adding columns can only decrease the relative volume and increase the condition number.
  2. 2.
    It therefore suffices to show the lack of a PTAS by establishing approximability thresholds for matrices with 2 columns.
  3. 3.
    Monotonicity implies that matrices with more columns cannot cross the approximability thresholds.
    This shows that the following criteria do not admit a PTAS: maximization of relative volume (Theorem 2), volume (Theorem 4) and S-optimality (Theorem 5); two-norm minimization (Theorem 7); pseudo-inverse minimization in the two- (Theorem 9), Frobenius (Theorem 10) and Schatten $p$-norms (Theorem 11); condition number minimization in the two- (Theorem 13), Frobenius (Theorem 14) and mixed norms (Theorem 15); and maximization of the stable rank (Theorem 17).
@@ 1.5 Literature
We review existing work on column subset selection, focused on computational complexity. Early surveys of column subset selection algorithms include [8, 25].
@@
NP hardness and PTAS inapproximability are shown in [9, Theorems 4 and 7] for the two-norm. Randomized and deterministic algorithms for wide matrices $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with $m<n$ are given in [1, 2, 29, 34, 35, 37].
In the context of A-optimal design, trace minimization corresponds to minimization of the pseudo-inverse in the Frobenius norm [34, 35]. For instance, if $A$ has full row rank then $\boldsymbol{A}^{\dagger}=\boldsymbol{A}^{T}(\boldsymbol{A}\boldsymbol{A}^{T})^{-1}$ and
$\|\boldsymbol{A}^{\dagger}\|_{F}^{2}=\operatorname{\mathrm{trace}}\left((\boldsymbol{A}^{\dagger})^{T}\boldsymbol{A}^{\dagger}\right)=\operatorname{\mathrm{trace}}\left((\boldsymbol{A}\boldsymbol{A}^{T})^{-1}\right).$
The reduction for $m=k$ is based on Odd Cycle Packing, which consists of finding a maximal family of vertex-disjoint odd cycles in a simple undirected graph.
@@
NP hardness and PTAS inapproximability are shown in [9, Theorems 4 and 6], while the exponential inapproximability of the volume is presented in [10, Theorem 5], [18, Theorem 8]. The latter involves a reduction based on Odd Cycle Packing for matrices $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{A})=m$, $k=m$, $m=1320d$ and $n=2112d$ for some $d>0$.
An early result [42, Theorem 3] in the context of D-optimality shows the NP-hardness of volume maximization via a reduction from Hamiltonian Circuit, where $k\geq n$ rows $\boldsymbol{C}\in\mathbb{R}^{k\times n}$ from an $m\times n$ adjacency matrix are to be selected to maximize $\operatorname{\mathrm{vol}}(\boldsymbol{C})^{2}=\det(\boldsymbol{C}^{T}\boldsymbol{C})$.
Greedy algorithms are presented in [9, 14], and local deterministic and randomized algorithms in [22, 23, 33, 36]. This is extended to Hilbert spaces in [30]. The importance of local volume maximization for revealing the rank is demonstrated in [13].
The S-optimality criterion bears a resemblance to the relative volume and was introduced in [39, (3.5)] and [31, (3.10)].
@@
The NP completeness in the Frobenius norm is established in [40, Theorem 2.2], and PTAS inapproximability conditional on the Unique Games Conjecture being true [6, Theorem 1.1]. Algorithms for the two- and Frobenius norms are given in [5].
Early randomized algorithms, based on volume sampling are given in [5, 16, 17]. A statistical perspective is provided in [41], and it is shown that subset selection is equivalent to the method of principal variables.
@@ 1.6 Assumptions
Singular value interlacing [26, 7.3.P44] implies
@@
We denote by $\boldsymbol{I}_{n}\equiv\begin{bmatrix}\boldsymbol{e}_{1}&\cdots&\boldsymbol{e}_{n}\end{bmatrix}\in{\mathbb{R}}^{n\times n}$ the identity matrix, and by $\boldsymbol{e}_{j}\in\mathbb{R}^{n}$ its columns.